A woman with a mass m = 68.0 kg is sitting on the edge of a large rotating disk with a mass M = 105 kg and a radius R = 3.20 m. The disk and woman are rotating at a rate ω = 0.500 rev/sec. Determine the moment of inertia I for the disk and woman assuming that the woman acts like a point mass at the edge of the disk. Calculate the total angular momentum for the disk and woman. If the woman walks to the center of the disk what is the new revolution rate for the disk?
Answer #1

I = MR^2/2 + mR^2 = 105*(3.20)^2/2 + 68*(3.20)^2 = 1233.92 kg.m^2

angular momentum = Iw = 1233.92*3.14 = 3874.50 kg.m^2s^-1

I1w1 = I2w2

105*(3.20)^2/2*W =   3874.50

=> W = 7.20 rad/ sec^-1